The sum of twice Kenesha's age and 22 is equal to her mother's age. If Kenesha's mother is 46 years old, how old is Kenesha?
Answer by checkley79(3341)

Let X be Kenesha's age.
Let 46 be her mom's age.
2X+22=46
2x=46-22
2X=24
X=24/2
X=12 age of Kenesha.
PROOF:
2*12+22=46
24+22=46
46=46